### v1

@@ 2 Related Work
3D content editing [62, 10, 41] has experienced significant advancements, particularly in the modification [52, 60] of geometry and appearance based on textual and spatial guidance. Several recent methods have shown remarkable promise using various inputs such as text prompts [33], sketches [29], and reference images [1, 14, 15], as well as leveraging models like CLIP [43, 45, 17] and diffusion models [33, 16, 39, 57, 6, 18, 58, 59, 9, 51, 54] to provide editing directions. Among these works, diffusion-based methods, in particular, have significantly enhanced the effectiveness of 3D content editing by utilizing the strong capabilities of diffusion models in text-based generation and editing. However, these methods primarily focus on localized edits, altering existing objects in the 3D scene rather than creating new objects from scratch or removing objects. In recent years, there has been a growing body of work focusing on object removal in 3D scenes by exploiting techniques such as differentiable mesh optimization [44], multi-view segmentation guidance [55, 32], and depth priors [23]. These methods typically involve delineating the editing boundaries in different viewpoints based on the shape of the target object. Consequently, they are unable to generate new objects in a scene from scratch without new object-related priors [40]. In contrast, our work targets the creation of 3D-consistent, high-quality, and background-harmonious new objects in 3D scenes from scratch.
@@ 2.2 SDS-based Object Insertion
To facilitate object generation and insertion in 3D scenes, some studies [13, 8, 4] have integrated the classic Score Distillation Sampling (SDS) algorithm [34] into 3D generation. This integration aims to leverage SDS’s robust generative capabilities to directly produce viewpoint-consistent 3D content within scenes. Drawing upon prior knowledge from 2D diffusion, SDS optimizes the rendered images of 3D models to align with the desired image distribution conditioned on text. This ensures that the edited 3D model meets the desired outcomes across various viewpoints, enabling the generation and insertion of 3D content from scratch. ReplaceAnything3D [4] enhances this approach by incorporating an improved version of SDS, known as HiFA [61], for generative object insertion, thereby enhancing the realism of editing effects. MVIP-NeRF [8] proposes a multi-view SDS that concurrently supervises surface normals and scene appearance, providing stronger supervision for scene editing and achieving improved results. While SDS-based methods effectively achieve viewpoint-consistent 3D content generation and editing, the inherent stochasticity in their optimization process often results in noticeable quality issues, such as over-smoothness [22, 47] and high saturation levels [25, 11]. In contrast, our method avoids these issues by using SDS optimization solely to provide coarse initial priors rather than the final editing output. We then design a multi-view diffusion model to inpaint view-consistent editing results.
@@ 5.3 Editing Results
Furthermore, as depicted in Fig. 7, the editing results produced by our model maintain a high level of realism and viewpoint consistency across a broad range of observation angles. Even in scenarios involving complex shapes, such as the flower case, our method consistently generates realistic and harmonious 3D results across various viewing angles, whether close or distant, high or low, highlighting the robustness of our approach.
@@ 6 Conclusion
Limitations. While our model excels at generative object insertion, it does have certain limitations. One key challenge is the lack of publicly available sufficient training data, which limits the model’s ability to support 360-degree content generation, as illustrated in Fig. 11. This limitation could be overcome as more training data becomes available in the future. Another limitation is that the method cannot be directly extended to effective object removal, as shown in Fig. 12. The main challenge is that the newly patched background often fails to align seamlessly with the original scene, resulting in reconstruction artifacts. Implementing distinct inpainting strategies for void areas and background regions within the editing zone could help address this issue, and we plan to explore this approach in future work.

### v2

@@ 2 Related Work
3D content editing [66, 10, 44, 29] has experienced significant advancements, particularly in the modification [55, 63] of geometry and appearance based on textual and spatial guidance. Several recent methods have shown remarkable promise using various inputs such as text prompts [35], sketches [31], and reference images [1, 14, 15], as well as leveraging models like CLIP [46, 48, 17] and diffusion models [35, 16, 42, 60, 6, 18, 61, 62, 9, 54, 57] to provide editing directions. Among these works, diffusion-based methods, in particular, have significantly enhanced the effectiveness of 3D content editing by utilizing the strong capabilities of diffusion models in text-based generation and editing. To enhance the controllability of 3D editing, SketchDream [23] introduces a sketch-based multi-view image generation diffusion model, enabling sketch-based text-to-3D generation and editing in object-centric cases. On the other hand, QNeRF [36] suggests extracting query features of the self-attention layers in the diffusion model during multi-view editing, then training a neural field to consolidate these queries, thereby enhancing editing consistency. However, these methods either primarily focus on localized edits, altering existing objects in the 3D scene, or solely support scene-independent, object-centric 3D content generation. This poses challenges for creating new objects from scratch or removing objects within a given scene. In recent years, there has been a growing body of work focusing on object removal in 3D scenes by exploiting techniques such as differentiable mesh optimization [47], multi-view segmentation guidance [58, 34], and depth priors [24]. These methods typically involve delineating the editing boundaries in different viewpoints based on the shape of the target object. Consequently, they are unable to generate new objects in a scene from scratch without new object-related priors [43]. In contrast, our work targets the creation of 3D-consistent, high-quality, and background-harmonious new objects in 3D scenes from scratch.
@@ 2.2 SDS-based Object Insertion
To facilitate object generation and insertion in 3D scenes, some studies [13, 8, 4] have integrated the classic Score Distillation Sampling (SDS) algorithm [37] into 3D generation. This integration aims to leverage SDS’s robust generative capabilities to directly produce viewpoint-consistent 3D content within scenes. Drawing upon prior knowledge from 2D diffusion, SDS optimizes the rendered images of 3D models to align with the desired image distribution conditioned on text. This ensures that the edited 3D model meets the desired outcomes across various viewpoints, enabling the generation and insertion of 3D content from scratch. ReplaceAnything3D [4] enhances this approach by incorporating an improved version of SDS, known as HiFA [64], for generative object insertion, thereby enhancing the realism of editing effects. TIP-Editor [65] introduces a stepwise 2D personalization strategy and a 3D scene editing framework incorporating SDS loss. This system enables 3D editing, such as object insertion, guided not only by the text prompt but also by a reference image. MVIP-NeRF [8] proposes a multi-view SDS that concurrently supervises surface normals and scene appearance, providing stronger supervision for scene editing and achieving improved results. While SDS-based methods effectively achieve viewpoint-consistent 3D content generation and editing, the inherent stochasticity in their optimization process often results in noticeable quality issues, such as over-smoothness [22, 50] and high saturation levels [26, 11]. In contrast, our method avoids these issues by using SDS optimization solely to provide coarse initial priors rather than the final editing output. We then design a multi-view diffusion model to inpaint view-consistent editing results.
@@ 5.1 Experimental Setup
Figure 8: Generation of Reflective Objects. Our model also supports the generation of reflective objects. The generated surfaces exhibit prominent view-dependent effects, such as specular reflections.
@@ 5.2 Comparisons
In addition to outperforming in qualitative comparison, our method also stands out in quantitative evaluation. As shown in Table 1, our method generates effective editing that best align with the text prompts across multiple viewpoints, achieving top scores in the CLIP-relevant metrics CTIS and DTIS. Additionally, the objects generated by our method exhibit a highly realistic appearance and blend seamlessly with the scene background, leading to superior performance in the MUSIQ metric that assesses the authenticity of rendered images.
@@ 5.3 Editing Results
Furthermore, as depicted in Fig. 7, the editing results produced by our model maintain a high level of realism and viewpoint consistency across a broad range of views covered by our inpainting. This highlights that within this coverage, the reconstructed 3D objects consistently display well-reconstructed appearances when viewed from arbitrary angles. Even in scenarios involving complex shapes, such as the flower case, our method consistently generates realistic and harmonious 3D results across various viewing angles, whether close or distant, high or low, highlighting the robustness of our approach.
Besides, as depicted in Fig. 8, our model also facilitates the generation of reflective objects. For instance, in the candlestick case in the top row, the central portion of the zoomed-in area does not have a highlight in the initial viewpoint. Upon transitioning to the second viewpoint, a leaf-shaped highlight becomes visible. Subsequently, in the third viewpoint, this highlight expands and blends with the surrounding highlights. This outcome showcases the efficacy of our model in producing objects with view-dependent effects.
@@ 6 Conclusion
Limitations. While our model excels at generative object insertion, it does have certain limitations. One key challenge is the lack of publicly available sufficient training data, which limits the model’s ability to support 360-degree content generation, as illustrated in Fig. 12. This limitation could be overcome as more training data becomes available in the future. Another limitation is that the method cannot be directly extended to effective object removal, as shown in Fig. 13. The main challenge is that the newly patched background often fails to align seamlessly with the original scene, resulting in reconstruction artifacts. Implementing distinct inpainting strategies for void areas and background regions within the editing zone could help address this issue, and we plan to explore this approach in future work. In addition, our current model has difficulty handling shadows from inserted objects. To ensure the newly-generated background remain unchanged, we extract the target object after inpainting and reintegrate it into the original background for subsequent reconstruction. However, as depicted in Fig. 14, the segmentation model sometimes overlooks shadows, resulting in reconstructions without them. Manual adjustments to the mask or using advanced segmentation methods could resolve this issue, as demonstrated by the manually segmented cases on the right side of Fig. 14.
@@
Figure 14: Limitation of Shadow Generation. To preserve the scene background, we segment the target object from the inpainted frames and reintegrate it into the original background for subsequent reconstruction. Nevertheless, the segmentation model may fail to account for shadows, resulting in reconstructions that lack this crucial element.
CRediT authorship contribution statement
Hongliang Zhong: Software, Data curation, Formal analysis, Validation, Investigation, Visualization, Methodology, Writing – original draft. Can Wang: Writing – review & editing. Jingbo Zhang: Writing – review & editing. Jing Liao: Conceptualization, Resources, Supervision, Funding acquisition, Project administration, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve the readability and language of the paper. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Ethical Approval
This study does not contain any studies with Human or animal subjects performed by any of the authors.
Acknowledgments
The work described in this paper was fully supported by a GRF grant from the Research Grants Council (RGC) of the Hong Kong Special Administrative Region, China [Project No. CityU 11208123].